Let 𝑔(πœƒ) = sin πœƒ πœƒ

a) Make a table of the values of g at values of u that approach πœƒ0 = 0 from above and below. Then estimatelim 𝑔(πœƒ)

πœƒβ†’0

b) Support your conclusion in part (a) by graphing g near πœƒ0 = 0


1
Expert's answer
2022-01-18T06:06:53-0500

a)

"g(\theta)=\dfrac{\sin \theta}{\theta}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \theta & \dfrac{\sin \theta}{\theta} \\ \hline
 -0.1 & 0.9983341665 \\
 \hline
 -0.01 & 0.9999833334 \\
 \hline
 -0.001 & 0.9999998333 \\
 \hline
 -0.0001 & 0.9999999983 \\
 \hline
 0.0001 & 0.9999999983 \\
 \hline
 0.001 & 0.9999998333 \\
 \hline
 0.01 & 0.9999833334 \\
 \hline
 0.1 & 0.9983341665 \\
 \hline

\end{array}"


"\lim\limits_{\theta\to 0}\dfrac{\sin \theta}{\theta}=1"

b)